OpenStudy (loser66):

Prove: If A is self-adjoint and if A^k=1 for some strictly positive integer k, then A^2=1 Please, help

OpenStudy (loser66):

I appreciate any tip, pleaaaaaaaaaase

OpenStudy (usukidoll):

:(

OpenStudy (loser66):

Please

OpenStudy (usukidoll):

depends on the subject. what is this?

OpenStudy (loser66):

theoretical linear algebra

OpenStudy (usukidoll):

don't know sorry D:

OpenStudy (loser66):

No worry, I will ask my prof for help although the chance to get help from him is low. hihihihi

OpenStudy (usukidoll):

my question is worse :(

OpenStudy (loser66):

what course? I know your level is higher than me, but still want to take a look

OpenStudy (usukidoll):

intro to advanced mathematics ... I'll bump my question

OpenStudy (loser66):

Woah.... above my head, hihihihi

OpenStudy (loser66):

@KingGeorge any tip?

OpenStudy (kinggeorge):

What are some of the properties of self-adjoint matrices that you're allowed to use?

OpenStudy (loser66):

sorry, I confuse , that is A* =A

OpenStudy (loser66):

e. values are real

OpenStudy (kinggeorge):

Hm. I'm not completely sure. I'll have to think about it some.

OpenStudy (loser66):

Thanks in advance. I need it for my final review. If you have something, please post it.

OpenStudy (kinggeorge):

Well what if we let \(\vec{v}\) be an eigenvector of \(A\) with eigenvalue \(\lambda\). Then\[A^k\vec{v}=\lambda^k\vec{v}\]and\[A^k\vec{v}=I\vec{v}=\vec{v}\]So \(\lambda^k=1\) which implies that \(\lambda=\pm1\). I'm not sure how to continue from here, but http://en.wikipedia.org/wiki/Unitary_matrix claims that what we have is equivalent to \(A\) being unitary, and therefore \[AA^*=I\implies A^2=I\]